OpenStudy (anonymous):

Find the linearization L(x) of f(x) at x=a. F(x)=x+(1/x) a=1

OpenStudy (anonymous):

I have done it but I do no have the right answer.

OpenStudy (solomonzelman):

your function is, \(\Large\color{black}{F(x)=\frac{x+1}{x} }\) , and ` a = 1 `. Correct?

OpenStudy (solomonzelman):

\(\large\color{black}{\frac{d}{dx} (\frac{x+1}{x})=\frac{(x+1)'~x~-~x'(x+1)}{x^2} }\)

OpenStudy (solomonzelman):

you just need the derivative of 1) ` x + 1 ` : 2) ` x ` :

OpenStudy (anonymous):

so 1+0=1 and 1

OpenStudy (solomonzelman):

yes, the derivative of both of the functions, of x+1, and of x, is 1.

OpenStudy (solomonzelman):

So, now fund the derivative of the (x+1)/x based on the aforementioned quotient rule.

OpenStudy (solomonzelman):